OpenStudy (anonymous):

HELP IMMEDIATELY PLEASE!!! Jimmy Bullard and his friend went for fishing. They got six fishes without heads, nine fishes without the tail and eight fishes cut in two halves. How many fishes did they caught ?

OpenStudy (jennyrlz):

Add the fishes ^.^

OpenStudy (anonymous):

how! please help

OpenStudy (anonymous):

theres six without heads so we have six so far,

OpenStudy (jennyrlz):

they try to throw you off by giving you extra info

OpenStudy (anonymous):

nine fishes and eight more,

OpenStudy (anonymous):

so add 6+9+8=

OpenStudy (anonymous):

so in total 0?

OpenStudy (jennyrlz):

so 6 + 9 + 8

OpenStudy (anonymous):

I think 0 because none of the fish were alive

OpenStudy (anonymous):

that doesn't matter they are still fish

OpenStudy (jennyrlz):

It doesnt ask for how many live fishes :P

OpenStudy (anonymous):

ok

OpenStudy (jennyrlz):

fish still smell like fish dead or alive xD

OpenStudy (anonymous):

so add 6+9+8= your answer

OpenStudy (jennyrlz):

^

OpenStudy (anonymous):

LOLOLOLOL CYA!

OpenStudy (jennyrlz):

Q.Q

OpenStudy (anonymous):

thats not the correct answer

OpenStudy (anonymous):

can you help we with another one T_T im about to cry

OpenStudy (anonymous):

please!

OpenStudy (anonymous):

add 6+9+8 thats the answer for your first question

OpenStudy (anonymous):

Take 9 from 6, 10 from 9, 50 from 40 and leave 6. How Come ??

OpenStudy (anonymous):

ok are you working with negative numbers?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

ok so here is the trick with negative numbers

OpenStudy (anonymous):

subracts 6 - 9 and you get three, then just add a - to the three and thats your answer

OpenStudy (anonymous):

WAIT I GET IT!!

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

SIX – 9 (IX) = S 9 (IX) – 10 (X) = I 40 (XL) – 50 (L) = X ===>SIX

OpenStudy (anonymous):

yeah you got it!

OpenStudy (anonymous):

yay!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I FIGURED OUT THE SOLUTION TO LIFE!! ITS 6!!!

OpenStudy (anonymous):

hahaha yeah

OpenStudy (anonymous):

did i trix you

OpenStudy (anonymous):

what

OpenStudy (anonymous):

A bus driver was heading down a street in Delhi. He went right past a stop sign without stopping, he turned left where there was a 'no left turn' sign and he went the wrong way on a one-way street. Then he went on the left side of the road past a cop car. Still - he didn't break any traffic laws. Why not?

OpenStudy (anonymous):

yes that would be right